OpenStudy (pixiedust1):

The period p of a pendulum, or the time it takes the pendulum to make one complete swing, varies directly as the square root of the length L of the pendu- lum. If the period of a pendulum is 1.5 s when the length is 2 ft, find the period when the length is 5 ft. Round to the nearest hundredth

OpenStudy (vickyisthesmartone123):

What do you think the answer is? :) State your process.

OpenStudy (raffle_snaffle):

@Vickyisthesmartone123 at least show OP how to start the problem.

OpenStudy (pixiedust1):

Okay, so I was thinking that 2/1.5=1.33

OpenStudy (pixiedust1):

1.33*5=6.7

OpenStudy (phi):

have you studied "direct variation" ?

OpenStudy (pixiedust1):

Not familiar with the tem

OpenStudy (pixiedust1):

*term

OpenStudy (phi):

I suspect you (should!) have learned it. the simple equation if y directly varies with x is y = k x where k is some constant (which in these problems you have to first find) your problem is a bit more complicated, because p varies directly with the square root of L which , as an equation means \[ p= k \sqrt{L} \]

OpenStudy (pixiedust1):

eeek that sounds complicated o_O

OpenStudy (phi):

the steps are to use the first part of the info If the period of a pendulum is 1.5 s when the length is 2 ft to find what "k' is then once you know k, you use it and the other info the length is 5 ft. to find the period

OpenStudy (phi):

you can give it a try. you start with \[ p = k \sqrt{L}\] now look at the info the period of a pendulum is 1.5 s when the length is 2 ft what number should we use for p and what number for L ?

OpenStudy (pixiedust1):

The p is 1.5 and the L is 2?

OpenStudy (phi):

yes. if you use those numbers instead of the letters in the "variation equation" \[ p = k \sqrt{L} \] what do you get ? i.e. what is the equation with numbers instead of letters?

OpenStudy (pixiedust1):

1.5 = k sqroot of 2

OpenStudy (phi):

do you know how to "solve for k" ?

OpenStudy (phi):

you should divide both sides by sqr(2)

OpenStudy (pixiedust1):

Hmm... no idea, but I think k=1.06?

OpenStudy (pixiedust1):

Aww darn I just found out that my mom says I gtg, I am so so sorry, but than you for helping me set up the problem, I will try to figure out the rest of it later. thank you so much for teaching me the concept up to this point!!

OpenStudy (phi):

\[ \sqrt{2}k= 1.5\\ \frac{\sqrt{2}}{\sqrt{2}}k= \frac{1.5}{\sqrt{2}} \\ k= \frac{1.5}{\sqrt{2}} \] yes k=1.06, but because they want the answer to the nearest 100th I would keep more digits k= 1.06066

OpenStudy (pixiedust1):

Oh okay cool, wait so 1.06 is the final answer? that's all the steps to solve the problem?

OpenStudy (phi):

now you put in 1.06066 for k and 5 for L and do the math, to find p

OpenStudy (pixiedust1):

Alright, thank you Phi!! and I am sorry for cutting it short, will use the concept to figure it out later!